Circular No. 6850

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1998Y IN NGC 2415 W. Li, M. Modjaz, R. R. Treffers, and A. V. Filippenko, University

of California at Berkeley, report their discovery of a supernova during the course of the Lick Observatory Supernova Search (cf. IAUC 6627) with the 0.8-m Katzman Automatic Imaging Telescope. SN 1998Y was found and confirmed on unfiltered CCD images obtained on Mar. 16.2 (mag ∼ 18.3) and 18.2 UT (mag ∼ 18.4), and it is located at α = 7^{h}36^{m}56^{s}.93, δ = +35^{o}14′36′′.0 (equinox 2000.0), which is ∼ 7′′ east and 6′′ north of the nucleus of NGC 2415. An image of the same field taken on 1997 Dec. 22 shows no star at this position. A spectrum (range 540–920 nm) obtained on Mar. 26 with the Keck-2 telescope by Filippenko, D. C. Leonard, and A. G. Riess (also of Berkeley) shows that SN 1998Y is of type II; the prominent Hα line has a P-Cyg profile.

SUPERNOVA 1998Z IN ANONYMOUS GALAXY L. Germany and the Mount Stromlo Abell Cluster Supernova Search

Team (cf. IAUC 6639, 6820) reports the discovery of an apparent supernova (V = 19.6) on V and R CCD images taken on Mar. 16 (limiting mag V = 20.5) by S. Sabine at the Mount Stromlo 1.27-m telescope (+ Macho Camera). SN 1998Z is located near Abell 3341 at α = 5^{h}24^{m}58^{s}.78, δ = −31^{o}21′38′′.1 (equinox 2000.0), which is 1′′.5 west and 2′′.3 south of the center of the host galaxy. SN 1998Z was not visible on similar images taken on Feb. 18 (limiting mag V = 20.5), but it is present on B, V, R, and I images taken at the Mount Stromlo 1.88-m telescope on Mar. 22 (limiting mag V = 22). A nearby star has position end figures 55^{s}.19, 20′19′′.8.

SUPERNOVA 1998aa IN ANONYMOUS GALAXY J. Maza, Department of Astronomy, University of Chile, reports the

discovery by Roberto Antezana of a supernova (B ∼ 18) on a film (T-Max 400) taken by L. E. Gonzalez using the Maksutov telescope at Cerro El Roble on Feb. 26.201 UT. SN 1998aa is located at α = 12^{h}25^{m}32^{s}.65, δ = +7^{o}38′59′′.0 (equinox 2000.0), which is 10′′.4 west and 0′′.8 south of the nucleus of a spiral galaxy. The object was confirmed by L. Gonzalez at Las Campanas using the 1-m Swope telescope on Mar. 9. On Mar. 24, a spectrum of the object was obtained by M. T. Ruiz (Department of Astronomy, University of Chile) using the 3.6-m telescope at the European Southern Observatory, from which M. Phillips (Cerro Tololo) classified the spectrum as a normal type-Ia supernova, 35 days after maximum.

1998 March 26  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green